OpenStudy (anonymous):

need help (-7b5 + 42b4) ÷ 7b

OpenStudy (anonymous):

is this factoring

OpenStudy (anonymous):

dont know it is in lib.arts math

OpenStudy (anonymous):

ok hol up

OpenStudy (mathmate):

When a polynomial is divided by a monomial, divide term by term. Example: \(\dfrac{3x^4+6x^3+9x^2}{3x}=\dfrac{3x^4}{3x}+\dfrac{6x^3}{3x}+\dfrac{9x^2}{3x}=x^3+2x^2+3x\)